OpenStudy (anonymous):

Solve the equation. 3(2x - 1)=1/2(4x - 2) + 2

OpenStudy (anonymous):

First start distributing things into the brackets, ending up with: 6x-3=2x-1+2 Then, collect similar terms - ie, the x terms and the non-x terms: 6x-2x=3-1+2 Then just add and subtract: 4x=4 and then divide both sides by 4 to get x by itself, and you have the answer: x=1

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Best response? :)